# Particles before symmetry

### v4

@@ 4.2 The VB-POV presentation of the Yukawa mechanism
whereas vector bosons are replaced by the corresponding affine covariant derivatives, e.g. $\nabla^{1},\nabla^{2},\nabla^{3}$ (see \@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025a)\@BBCP for more details).
@@
In this formulation, weak isospin $T_{3}$—defined only with respect to a chosen basis of the Lie algebra—has no independent geometrical meaning (see B.2 and footnote 41). Accordingly, left-handed fermions, together with $\varphi^{+}$ and $\varphi^{0}$ (also $SU(2)$-doublets), are best understood simply as components of the vector fields $\mathbf{Q}_{L}$ and $\mathbf{L}_{L}$ (where I have introduced boldface to emphasise their vectorial character). The familiar distinction between, say, the electron and the electron-neutrino, or between the up- and down-left quarks, does not arise at this level: it appears only through their couplings with the Higgs. The Higgs field $\boldsymbol{\varphi}$ provides a frame within $\mathbb{C}^{2}$ that endows $T_{3}$—and hence these component fields—with physical significance. The charges listed in Table 1 are already adapted to this frame, since they presuppose the choice $\varphi=(0,\varphi^{0})^{T}$ (i.e. $\varphi^{+}=0$); for example, only in that frame do the left-handed up-quark components appear as $(u^{I}_{L},0)^{T}$. [^19] This explains why Table 1, reproduced from \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Table 8.2)\@BBCP, can be misleading: if both components of the Higgs are retained, the up and down components of the left-handed quarks and leptons do not yet have any physical meaning.
@@
Thus, geometrically, it makes more sense to define the left-handed components of both leptons and quarks as parallel and orthogonal to the Higgs according to the inner product on $E^{2}$, i.e.:
$\mathbf{e}_{L}$ $:=\langle\mathbf{L}_{L},e_{0}\rangle_{2}\,e_{0},\quad\text{with}\quad e_{L}=\langle\mathbf{L}_{L},e_{0}\rangle_{2}\,;\quad\boldsymbol{\nu}_{eL}:=\mathbf{L}_{L}-\boldsymbol{e}_{L},$ (4.11)
$\mathbf{u}^{I}_{L}$ $:=\langle\mathbf{Q}^{I}_{L},e_{0}\rangle_{2}\,e_{0}\quad\text{with}\quad u^{I}_{L}=\langle\mathbf{Q}^{I}_{L},e_{0}\rangle_{2}\,;\quad\mathbf{d}_{L}:=\mathbf{Q}_{L}-\mathbf{u}_{L},$ (4.12)
where capital $I$ indicates color components (i.e. red, green and blue; but note that $\mathbf{Q}^{I}_{L}$ is still a vector in $E^{2}$) in an orthonormal frame of $\mathbb{C}^{3}$ and I used the notation $e_{0}$ (without boldface, to avoid confusion with the electron) for the unit-direction of the Higgs, introduced in Section 3.1 (not to be confused with the left-handed electron, $\mathbf{e}_{L}$). [^20] Not many textbooks that I have encountered emphasise this point—\@BBOPcitep\@BAP\@BBN(Tong, 2025)\@BBCP is an exception. And none describe it geometrically as I did here.
Before we give the geometric interpretation of (4.8), and of the corresponding form for quarks, note that, given an orthonormal basis for $E^{2}$, we can form duals: for $\xi=\xi^{\perp}e_{\perp}+\xi^{\parallel}e_{0}=(\xi^{\perp},\xi^{\parallel})^{T}$ (e.g. $\mathbf{e}_{L}=\mathbf{L}_{L}^{\parallel},\,\boldsymbol{\nu}_{eL}=\mathbf{L}_{L}^{\perp}$) the dual takes the conjugate of the transpose, so:
$((\xi^{\perp},\xi^{\parallel})^{T})^{*}=(\overline{\xi}^{\perp},\overline{\xi}^{\parallel}).$ (4.13)
Moreover, given Hermitian bundles $E,F$, we use $\langle\cdot,\cdot\rangle_{E}$, $\langle\cdot,\cdot\rangle_{F}$ for fibrewise pairings, and write iterated contractions as $\langle\langle\,.\,,\,.\,\rangle_{E},\,.\,\rangle_{F},$ with obvious extensions to tensor products.
Now using (4.13) and an orthonormal frame aligned with the Higgs, the Yukawa term for the leptons in Equation (4.8) now can be stated directly using (trivialisation-independent) sections of the vector bundles, without the need to involve the sections $\sigma$ of the principal bundles, and reads (including a coupling constant, $g_{e}$):
@@
where the first equality gives the ‘standard’ definition; $e_{R}$ and $e_{L}$ are, right and left-handed (resp.) Weyl spinor but internal scalars ($e_{L}$ is the magnitude of $\mathbf{L}_{L}$ along the Higgs, as in (4.11)); $\langle.\,,\,.\rangle_{2}$ is complex anti-linear in the first entry acting on the relevant spaces as:
$\langle.\,,\,.\rangle_{2}:E^{2}\otimes(\otimes^{3}E^{1*})\times E^{2}\otimes(\otimes^{3}E^{1})\rightarrow(\otimes^{6}E^{1}).$ (4.15)
And $\langle.\,,\,.\rangle_{1}$ is just the scalar inner product in $\mathbb{C}$. [^21] This is slightly misleading: what we have here is that $\varphi\in\Gamma(E^{2}\otimes^{3}E^{1})$, i.e. the third tensor product of $E^{1}$, which is still one-dimensional, $e^{*}_{L}\in\Gamma(E^{2*}\otimes^{3}E^{1*})$, and $e_{R}\in\Gamma(\otimes^{6}E^{1})$. This is why they match to a scalar. From (4.14) we can see how mass terms, proportional to $g_{e}v$ (as well as interactions with the Higgs field) emerge for the electron.
Geometrically, the inner products in (4.14) are a very natural way to obtain scalars: we are measuring ‘internal angles’ between the different particles seen as vector fields on the same spaces. Thus I take this form of (4.14), namely $\langle\langle\mathbf{L}_{L},\boldsymbol{\varphi}\rangle_{2},e_{R}\rangle_{1}$, to be a more transparent interpretation of the Yukawa term for leptons. Once one determines the inner product, there is no geometric justification to scale each such term by a complex number (or, more generally, to find alternative maps involving the group centraliser).
@@
Chirality shows up in the fact that only left-handed particles have components in $E^{2}$; right-handed particles do not. $e_{R}$ couples to the $\mathbb{C}^{1}$ components of $\mathbf{L}_{L}$ and $\boldsymbol{\varphi}$. Neutrinos don’t acquire mass, but not because they are orthogonal to the Higgs—which they are—but because we have not included right-handed neutrinos in our particle content. Had we included right-handed neutrinos, they would acquire mass by coupling to the ‘symplectic dual’ of the neutrino; this is what happens in the case of quarks. [^22] Because of this feature, the Yukawa terms for leptons are diagonal in generations: these mass terms don’t mix, say electrons with muons and taus.
This ‘symplectic dual’, called $\varphi_{c}$ on Table 1, is obtained by recruiting another geometric structure that we can equip $\mathbb{C}^{2}$ with (besides the Hermitean inner product): an orientation. This implies we can use the totally anti-symmetric form, or the volume form, $\epsilon_{ab}$, as part of the geometrical structure. In other words, whereas the Higgs mechanism, described in Section 3, used the structure $(E^{2},M,\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2})$, here we extend that to the structure $(E^{2},M,\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2},\epsilon)$. [^23] Under $A\in U(2)$, $\epsilon_{ab}$ is taken to transform as $\epsilon_{ab}\mapsto\det{(A)}\epsilon_{ab}$. So $SU(2)$ preserves it. Moreover, since $AA^{\dagger}=\mathbb{1}$ for any $A\in U(n)$, we know that $\det{(A)}\det{(A^{\dagger})}=|\det{(A)}|=1$, so $\det{(A)}=e^{i\theta}$ denotes an orientation change the $\mathbb{C}^{n}$. Using $\epsilon_{ab}$ as a geometric datum then implies we have a fixed orientation, as well as an inner product, on $\mathbb{C}^{2}$.
@@
Now, besides the metric, we can use $\epsilon_{ab}$ and its inverse $\epsilon^{ab}$ to raise or lower indices. [^24] Indeed, in standard differential geometry, we can find a similar sort of operator acting on two dimensions: the Hodge star: which would take a basis $e_{0},e_{1}\mapsto-e_{1},e_{0}$, respectively, so its action on vectors can be written in this frame as a matrix operator: $*=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, which is of the same form as $\epsilon_{ab}$. Thus if we call the isomorphism $J:E^{2}\rightarrow E^{2*}$ which acts as $\xi\mapsto\langle\xi,\cdot\rangle$ we have:
@@
where we used, in abstract index notation, $h_{ab}$ as the inner product on $E^{2}$. Thus we call
$\boldsymbol{\varphi}_{c}:=C(\boldsymbol{\varphi});$ (4.18)
it can be seen as a measure on the ‘areas’ orthogonal to $\boldsymbol{\varphi}$, which is why I called it a ‘symplectic dual’.
Denoting the generation by an index $i=1,2,3$, we then have, for the total Yukawa coupling term for quarks: [^25] Note that here, unlike for the leptons and the left-handed quarks, the up and down right-handed quarks are genuinely different particles, since they have different components in $E^{1}$.
@@
The boldface on lowercase letters is used to indicate that these are vector fields. This is the geometric form of the ‘standard’ definition (cf. \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Lemma 8.8.4)\@BBCP); it is, in the VB-POV, what really counts.
Nonetheless, as often is the case in physics, we can extract more information by introducing a frame: here, once more it is convenient, in order to compare with standard presentations, to choose the orthonormal frame for the Higgs, which gives the components for the quarks along and orthogonal to the Higgs (given in Equation (4.11)) as in Table 1, as well as $\varphi^{+}=0$. Then:
@@
where now all variables are scalar (and we are summing over the color indices, $I$, as well as over the generations $i,j$).
The Yukawa matrices $Y$ are generically non-diagonal, i.e. they mix generations of quarks. One can always find linear combinations of quarks such that, say, $Y^{u}$ is diagonal; this defines what is called the mass basis. But $Y^{u}$ and $Y^{d}$ cannot be diagonalised simultaneously, and the residual mixing is encoded in the Cabibbo–Kobayashi–Maskawa (CKM) matrix. Most textbooks (cf. \@BBOPcite\@BAP\@BBN(Hamilton, 2017, p. 515)\@BBCP) then explain that the CKM matrix describes the physical effects of left-handed quark mixing across generations, from the ‘mass eigenstate basis’ to the ‘weak eigenstate basis’ (the latter being the one we have used here). It then “follows that the interactions with the W-bosons can connect quarks from different generations if the CKM matrix is not diagonal” (ibid).
@@
One might have thought that, from the geometry-first perspective, the three generations of quarks could not coherently mix. After all, if different generations correspond to sections associated with different masses, they might seem to inhabit distinct vector bundles, and sections of distinct bundles cannot be linearly combined. Yet this worry dissolves once we note that the generations do not correspond to different bundles at all, but to different sections of the same composite bundle. That is, all the three generations live in a direct sum, e.g.
$(E^{1}\otimes E^{2}\otimes E^{3})\oplus(E^{1}\otimes E^{2}\otimes E^{3})\oplus(E^{1}\otimes E^{2}\otimes E^{3}).$ (4.21)
@@
So each quark field is a section of a tensor product of the fundamental bundles, together with a trivial “generation” factor,
$\mathbf{Q}_{L}^{i}\in\Gamma(S_{L}\!\otimes\!E_{3}\!\otimes\!E_{2}\!\otimes\!E_{1}^{Y_{Q}}\!\otimes\!\mathbb{C}^{3}_{\text{gen}}),\qquad i=1,2,3,$
and likewise for $u_{R}^{i}$ and $d_{R}^{i}$. We can use the linear structure of the fundamental bundles to form linear combinations of the three generations of quark. The Yukawa couplings then act as endomorphisms on this generation factor,
$Y_{u},\,Y_{d}\in\Gamma(\mathrm{End}(\mathbb{C}^{3}_{\text{gen}})).$
Diagonalising $Y_{u}$ or $Y_{d}$ corresponds simply to choosing an orthonormal frame in $\mathbb{C}^{3}_{\text{gen}}$, i.e. passing to the mass basis. But since $Y_{u}$ and $Y_{d}$ cannot be diagonalised simultaneously, their relative rotation
$V_{\text{CKM}}=U_{u}^{\dagger}U_{d}$
appears as a unitary automorphism of the generation fibre. Thus, from the VB-POV, the Cabibbo–Kobayashi–Maskawa matrix is nothing but a geometric rotation within the trivial generation bundle, fully compatible with the underlying fibre structure.
@@
In the geometric perspective one aspect of the situation is more transparent. If the up and down left-handed quarks were truly independent particles—i.e. distinct fields rather than components of the same field (usually called a doublet) in $E^{2}$—we could diagonalise $Y_{u}$ and $Y_{d}$ separately. But because they are components of the same $E^{2}$-field, and because these components are coupled to different fields (e.g. $\varphi,\varphi_{c}$), we can’t. Correspondingly, the $W$ bosons represent $\nabla^{2}$, the covariant derivative on $E^{2}$, and so they, too, naturally mix generations when they couple to the relevant currents. [^26] But other aspects of the Yukawa mechanism remain, to my eyes, mysterious from the VB-POV. For instance, it is a little disappointing that, unlike their left-handed counterparts, up and down right-handed quarks can’t be straightforwardly understood as components of a single vector field, due to their different components in $\mathbb{C}^{1}$. If they could be so understood, in place of (4.19), we would have the simpler: $T(\mathbf{Q}_{L},\boldsymbol{\varphi},\mathbf{Q}_{R})=Y^{d}_{ij}\langle\langle\langle\mathbf{Q}_{L}^{i},\mathbf{Q}_{R}^{j}\rangle_{3},\boldsymbol{\varphi}\rangle_{2}\rangle_{1}+Y^{u}_{ij}\langle\langle\langle\mathbf{Q}_{L}^{i},\mathbf{Q}_{R}^{j}\rangle_{3},\boldsymbol{\varphi}_{c}\rangle_{2}\rangle_{1},$ (4.22) which only takes the components of the same inner product along and orthogonal to the Higgs. But the obstruction is the hypercharge split between $u_{R}$ and $d_{R}$; it is this assignment, not the VB-POV machinery, that prevents a single right-handed multiplet with a unified coupling.
5 A defense of the geometry-first formulation
To motivate the methodological defense of the VB-POV, let me recall an illustrative example from the previous Section. The Yukawa coupling for quarks depends essentially on the orientation of $\mathbb{C}^{2}$: it requires the introduction of $\boldsymbol{\phi}_{c}$, which encodes the oriented area orthogonal to the Higgs. This shows how, in the VB-POV, the reduction from $U(2)$ to $SU(2)$ must arise directly from subvening geometric structures. Again the spacetime analogy is instructive. The automorphism group of the tangent bundle equipped with a Lorentzian metric is $SO(3,1)$ (or $O(3,1)$), a fact that becomes explicit once orthonormal frames are introduced. Yet most of spacetime geometry can be developed without ever invoking $SO(3,1)$; this is precisely why it serves as an analogue for the geometry-first VB-POV of gauge theory. And in the spacetime case, if one were to posit some other group acting on TMTM—say $O(2)$ or SU(n)SU(n)—a clear geometric rationale would be required: which feature of $TM$ could prompt one to consider such an action?
The broader moral is already clear: the VB-POV ties the existence of symmetry groups much more tightly to the underlying geometry of the vector bundle than the symmetry-first formulation requires. And while this may look like a liability of the VB-POV, it in fact strengthens its case over the PFB-POV.
This Section develops that defense in four stages. First, in Section 5.1, I show how the symmetry-first formulation, represented by the PFB-POV, permits a looseness between geometry and symmetry—a slack that can manifest even in perfectly ordinary gauge-theoretic settings. I analyse the conditions under which symmetry groups fail to reflect the geometry of the fibres and illustrate this failure through three representative examples that typify the range of possibilities allowed by the PFB-POV. Second, in Section 5.2, I exhibit two cases of groups for which the VB-POV could not possibly recover the full flexibility of the PFB-POV. One of these cases shows that quantisation of charge is more fundamental in the VB-POV. Third, in Section 5.3, I turn to composite fibres such as those appearing in the Standard Model, where different factors of the gauge group act on different tensor components—sometimes trivially. Here the slack of the PFB-POV becomes manifest: the symmetry group can no longer be recovered from the geometry of any single associated bundle. The VB-POV restores coherence by grounding each gauge factor in the automorphism group of its own fundamental vector bundle. Finally, in Section 5.4, I turn to the methodological upshot: the VB-POV’s refusal to countenance such slack is not a constraint to be lamented but a virtue to be embraced, since it aligns the theory’s ontology with its genuine explanatory structure.

### v5

@@ 4.2 The geometry-first presentation
whereas vector bosons are replaced by the corresponding affine covariant derivatives, e.g. $\nabla^{1},\nabla^{2},\nabla^{3}$ (see \@BBOPcitep\@BAP\@BBN(Gomes, 2024, 2025b)\@BBCP for more details).
@@
In this formulation, weak isospin $T_{3}$—defined only relative to a chosen basis of the Lie algebra—has no independent geometrical meaning (see Section B.2 and footnote 40). Left-handed fermions are best understood as components of vector fields $\mathbf{Q}_{L}$ and $\mathbf{L}_{L}$. The distinction between electron and electron-neutrino, or between up- and down-left quarks, appears only through their couplings to the Higgs. The Higgs field $\boldsymbol{\varphi}$ provides a frame in $\mathbb{C}^{2}$ that gives $T_{3}$ physical significance. The charges in Table 1 presuppose the frame $\varphi=(0,\varphi^{0})^{T}$; only in that frame do left-handed up-quark components appear as $(u^{I}_{L},0)^{T}$. [^21] This is why Table 1 can mislead: if both Higgs components are retained, the ‘up’ and ‘down’ labels have no physical meaning yet.
@@
Geometrically, one should define left-handed components as parallel and orthogonal to the Higgs:
$\mathbf{e}_{L}$ $:=\langle\mathbf{L}_{L},e_{0}\rangle_{2}\,e_{0},\quad e_{L}:=\langle\mathbf{L}_{L},e_{0}\rangle_{2}\,,\quad\boldsymbol{\nu}_{eL}:=\mathbf{L}_{L}-\boldsymbol{e}_{L},$ (4.11)
$\mathbf{u}^{I}_{L}$ $:=\langle\mathbf{Q}^{I}_{L},e_{0}\rangle_{2}\,e_{0},\quad u^{I}_{L}:=\langle\mathbf{Q}^{I}_{L},e_{0}\rangle_{2}\,,\quad\mathbf{d}_{L}:=\mathbf{Q}_{L}-\mathbf{u}_{L},$ (4.12)
where $I$ indexes colour (red, green, blue) and $e_{0}$ is the unit Higgs direction from Section 3.1. Thus left-handed up and down quarks are not distinct particles but components of the same field. [^22] Some textbooks note this—cf. \@BBOPcite\@BAP\@BBN(Tong, 2025, p. 185)\@BBCP—but usually in the context of unification ‘before symmetry breaking’. Here, these just are parallel and orthogonal components of the same field relative to the Higgs; treating them as independent particle fields would be misleading.
Given an orthonormal basis for $E^{2}$ aligned with the Higgs, duals are formed by conjugate transpose: for $\xi=(\xi^{\perp},\xi^{\parallel})^{T}$, we have $\xi^{*}=(\overline{\xi}^{\perp},\overline{\xi}^{\parallel})$. For Hermitian bundles $E,F$, we write iterated contractions as $\langle\langle\cdot,\cdot\rangle_{E},\cdot\rangle_{F}$.
With an orthonormal frame aligned to the Higgs, the lepton Yukawa term (4.8) becomes (with coupling $g_{e}$):
@@
where $e_{R}$ and $e_{L}$ are Weyl spinors that are internal scalars ($e_{L}$ is the magnitude of $\mathbf{L}_{L}$ along the Higgs). The inner product $\langle\cdot,\cdot\rangle_{2}$ is complex anti-linear in its first entry. [^23] More precisely: $\varphi\in\Gamma(E^{2}\otimes^{3}E^{1})$, $e^{*}_{L}\in\Gamma(E^{2*}\otimes^{3}E^{1*})$, and $e_{R}\in\Gamma(\otimes^{6}E^{1})$; these match to a scalar. (The $E^{3}$, $E^{2}$, and $E^{1}$ factors contract to $\underline{\mathbb{C}}$ via the fibre inner products; see Table 1 for the $E^{1}$-exponents.) Mass terms proportional to $g_{e}v$ emerge for the electron.
The inner products in (4.13) are geometrically natural: they measure internal angles between particle fields on the same spaces. Once the inner product is fixed, there is no geometric justification for scaling by complex numbers or invoking the group centraliser. The inner product does the real work.
@@
Chirality shows up here: only left-handed particles have components in $E^{2}$. Neutrinos do not acquire mass—not because they are orthogonal to the Higgs (which they are), but because we have not included right-handed neutrinos. Had we included them, they would couple to what I call the ‘symplectic dual’ of the Higgs; this is what happens for quarks. [^24] For this reason, lepton Yukawa terms are diagonal in generations: they do not mix electrons with muons or taus.
The symplectic dual $\varphi_{c}$ (Table 1) recruits a further geometric structure on $\mathbb{C}^{2}$: an orientation, encoded by the volume form $\epsilon_{ab}$. Whereas the Higgs mechanism used $(E^{2},M,\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2})$, here we extend to $(E^{2},M,\mathbb{C}^{2},\langle\cdot,\cdot\rangle_{2},\epsilon)$. [^25] Under $A\in U(2)$, $\epsilon_{ab}\mapsto\det(A)\epsilon_{ab}$, so $SU(2)$ preserves it. Since $|\det(A)|=1$ for unitary $A$, $\det(A)=e^{i\theta}$ encodes orientation. Fixing $\epsilon_{ab}$ as geometric data fixes both inner product and orientation on $\mathbb{C}^{2}$.
@@
Besides the metric, we can use $\epsilon_{ab}$ and $\epsilon^{ab}$ to raise or lower indices. [^26] Compare the Hodge star in two dimensions: $e_{0},e_{1}\mapsto-e_{1},e_{0}$, which acts as the matrix $\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, the same form as $\epsilon_{ab}$. Writing $J:E^{2}\rightarrow E^{2*}$ for the conjugate-linear isomorphism $\xi\mapsto\langle\xi,\cdot\rangle$ (conjugate-linear because the inner product is sesquilinear), we define
@@
where $h_{ab}$ is the inner product. The map $C$ is itself conjugate-linear, and so $\boldsymbol{\varphi}_{c}:=C(\boldsymbol{\varphi})$ measures areas orthogonal to $\boldsymbol{\varphi}$—hence ‘symplectic dual’.
With $i=1,2,3$ indexing generations, the quark Yukawa term is: [^27] Unlike leptons and left-handed quarks, right-handed up and down quarks are genuinely distinct particles, having different $E^{1}$ components.
@@
This is the geometric form of the standard definition (cf. \@BBOPcite\@BAP\@BBN(Hamilton, 2017, Lemma 8.8.4)\@BBCP). In the Higgs-aligned frame with $\varphi^{+}=0$:
@@
summing over colour $I$ and generations $i,j$.
The Yukawa matrices $Y$ mix generations. One can diagonalise $Y^{u}$ by passing to the mass basis, but $Y^{u}$ and $Y^{d}$ cannot be diagonalised simultaneously; their mismatch is encoded in the CKM matrix.
@@
One might worry that, from the geometry-first perspective, different generations cannot coherently mix—after all, if they correspond to different masses, they might inhabit distinct bundles. But this worry dissolves: generations are not different bundles but different sections of the same composite bundle. All three live in
$(E^{1}\otimes E^{2}\otimes E^{3})\oplus(E^{1}\otimes E^{2}\otimes E^{3})\oplus(E^{1}\otimes E^{2}\otimes E^{3}),$ (4.18)
@@
so each quark field is a section of a tensor product together with a trivial generation factor:
$\mathbf{Q}_{L}^{i}\in\Gamma(S_{L}\!\otimes\!E_{3}\!\otimes\!E_{2}\!\otimes\!E_{1}^{Y_{Q}}\!\otimes\!\mathbb{C}^{3}_{\text{gen}}),\qquad i=1,2,3.$
The Yukawa couplings act as endomorphisms on this factor, $Y_{u},Y_{d}\in\Gamma(\mathrm{End}(\mathbb{C}^{3}_{\text{gen}}))$. Diagonalising either corresponds to choosing a frame in $\mathbb{C}^{3}_{\text{gen}}$. Since both cannot be diagonalised simultaneously, their relative rotation $V_{\text{CKM}}=U_{u}^{\dagger}U_{d}$ appears as a unitary automorphism of the generation fibre. The CKM matrix is nothing but a geometric rotation within this trivial bundle.
@@
The geometric perspective clarifies one aspect. If up and down left-handed quarks were independent particles rather than components of a single $E^{2}$-field, we could diagonalise $Y_{u}$ and $Y_{d}$ separately. Because they are components of the same field, coupled to $\varphi$ and $\varphi_{c}$ respectively, we cannot. The $W$ bosons represent $\nabla^{2}$, the covariant derivative on $E^{2}$, and so they too naturally mix generations. [^28] Other aspects remain mysterious. Unlike their left-handed counterparts, right-handed up and down quarks cannot be understood as components of a single field, due to their different $E^{1}$ components. If they could, (4.16) would simplify to inner products along and orthogonal to the Higgs. The obstruction is the hypercharge split between $u_{R}$ and $d_{R}$; this assignment, not the geometry-first machinery, prevents a unified coupling.
5 A defence of the geometry-first formulation
The geometry-first picture binds symmetry to geometry tightly. One might take this for a defect. I shall argue that it is the picture’s chief recommendation.
In Section 1 I claimed that the principal-bundle picture tolerates a slack between symmetry and geometry that the geometry-first picture does not. It is time to make good on that claim. The slack is this: in the PFB-POV, the group $G$, the representation $\rho$, and the fibre $V$ can be chosen independently, subject only to mutual consistency. The geometry-first picture eliminates this freedom. [^29] Throughout this section, $\mathsf{Aut}(V)$ means the group of automorphisms preserving whatever fibre structure is in play (e.g. Hermitian form, and where specified a unit volume form).
I proceed in four steps. First I exhibit the slack through concrete examples (Section 5.1). Then I show that the geometry-first picture has genuine limitations: certain groups—notably the exceptional families—fall outside its natural reach (Section 5.2). Third, I show that the matter content of the Standard Model cannot, by itself, determine the gauge group—but that the VB-POV sidesteps this failure by starting from fundamental bundles (Sections 5.3). Finally, I argue that physicists already tacitly assume the VB-POV’s constraints (Section 5.4).